OpenStudy (anonymous):

i am to multiply. and leave my answer in exponential form for..(5y^2)(3y) please help

OpenStudy (anonymous):

wow

OpenStudy (lgbasallote):

hint \[\LARGE y^m \times y^n = y^{m+n}\]

OpenStudy (anonymous):

all greek

OpenStudy (lgbasallote):

let me do a demonstration

OpenStudy (anonymous):

those are placeholders/

OpenStudy (anonymous):

please

OpenStudy (anonymous):

what is that

OpenStudy (lgbasallote):

\[\Large (3x^2)(4x^3) \implies (3\times 4)(x^{2+3}) \implies 12x^5\] still greek?

OpenStudy (lgbasallote):

here's another demo \[\Large (5a)(4a^2) \implies (5\times 4)(a^{1+2}) \implies 20a^3\]

OpenStudy (anonymous):

where does the 4 come from

OpenStudy (lgbasallote):

that is an example...

OpenStudy (anonymous):

they are coefficient that can be any number depending on what was given

OpenStudy (lgbasallote):

\[\Large (2b)(3b) \implies (2\times 3)(b^{1+1}) \implies 6b^2\] there..no 4

OpenStudy (anonymous):

now i'm worse confused

OpenStudy (anonymous):

i want to cry

OpenStudy (lgbasallote):

haha where are you confused?

OpenStudy (lgbasallote):

b can be written as b^1 that's why there's a 1

OpenStudy (anonymous):

all of it..maybe its just brain farts

OpenStudy (anonymous):

where does the b come from

OpenStudy (lgbasallote):

those are examples

OpenStudy (lgbasallote):

im demonstrating how to multiply

OpenStudy (anonymous):

I think he needs the basic of all basics on exponents, igba

OpenStudy (anonymous):

here legs, http://www.purplemath.com/modules/exponent.htm

OpenStudy (lgbasallote):

i tried at the start...it was greek...

OpenStudy (lgbasallote):

that's why im using numbers now...

OpenStudy (anonymous):

i am female

OpenStudy (anonymous):

she, then

OpenStudy (anonymous):

it is still greek

OpenStudy (anonymous):

check out the weblink that I posted, legs

OpenStudy (anonymous):

where

OpenStudy (anonymous):

http://www.purplemath.com/modules/exponent.htm

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

enjoy :)

OpenStudy (anonymous):

thank you:(

OpenStudy (lgbasallote):

sad face? lol

OpenStudy (anonymous):

not lol...sad